OpenStudy (anonymous):

What is the percent (m/m) of a solution containing 25 g of NaCl in 175 g of solution? A. 700% B. 0.14% C. 7.0% D. 14% It isn't B

OpenStudy (anonymous):

why do u think it isnt B)????????????????

OpenStudy (anonymous):

well i feel the answer is B)

OpenStudy (anonymous):

I got the answer wrong and I chose B

OpenStudy (anonymous):

well, the answer IS 25/175=1/7=0.1428......

OpenStudy (anonymous):

ok...what is (m/m) according to u ?

OpenStudy (anonymous):

That's what I got as well, so I'm really confused.

OpenStudy (anonymous):

sorry ! really sorry ...answer is d)

OpenStudy (anonymous):

LOL i forgot to multiply by 100 ......truly sorry!!!!!!!!

OpenStudy (anonymous):

That makes much more sense. Thank you so much.